# Fine-tuning Reinforcement Learning Models is Secretly a Forgetting Mitigation Problem

### v2

@@ 2 Forgetting of pre-trained capabilities
In state coverage gap, we consider a pre-trained agent that is performing well mostly on Far and does not know how to behave on Close. However, when fine-tuned on Close, its behavior on Far will deteriorate considerably due to forgetting (Schaul et al., 2019) and will have to be re-acquired. This setting is representative of common transfer RL scenarios (Parisotto et al., 2015; Rusu et al., 2016, 2022), see also the top row of Figure 1 and Figure 2 for illustration.
@@ 3 Experimental setup
NetHack Learning Environment (Küttler et al., 2020) is a complex game consisting of procedurally generated multi-level dungeons. Since their layouts are randomly generated in each run, the player has to learn a general strategy rather than memorize solutions. NetHack is stochastic and requires mastering diverse skills, such as maze navigation, searching for food, fighting, and casting spells. It has been a popular video game for decades that recently has become a challenging testbed at the forefront of RL research (Hambro et al., 2022a; Piterbarg et al., 2023; Klissarov et al., 2023). Due to computational constraints, we focus solely on a single setting in our experiments, i.e., Human Monk.
@@ 5 Analysis: forgetting of pre-trained capabilities hinders RL fine-tuning
We study the extent to which knowledge retention techniques mitigate the negative effects of imperfect cloning gap on two levels representing Far states: level $4$ and Sokoban level [^1] In NetHack, the Sokoban level is a branch of the dungeon modeled and named after an NP-hard game where the goal is to push boxes on target locations, see NetHack wiki and Appendix B.1., see Figure 5. The performance of fine-tuning on level 4 can be temporarily enhanced by EWC and consistently improved by KS or BC, which is in line with the results presented in Figure 3(a). On the other hand, the Sokoban level poses an interesting challenge: solving it is not only an objective hard in its own right but also requires a skillset that is not immediately useful for the rest of the game.
This allows us to get some insights into the qualitative differences between the KS and BC. Namely, KS struggles with sustaining the performance on Sokoban, as uses trajectories gathered by the online policy. These do not contain any Sokoban levels at the start of the fine-tuning, as Sokoban is only encountered in the deeper parts of the dungeon. Conversely, BC uses data gathered by the expert and, as a result, constantly rehearses the correct way of solving this puzzle. As such, we note that both BC and KS have their specific advantages. We identify designing methods that combine these improvements as important future work. See Appendix D for additional NetHack metrics.
@@ 7 Limitations & Conclusions
8 Impact statement
@@ B.1 NetHack
We fine-tune the model pre-trained by Tuyls et al. (2023), which scales up (from 6M to 33M parameters) and modifies the solution proposed by the ’Chaotic Dwarven GPT-5’ team, which is based on Sample Factory (Petrenko et al., 2020) that was also used in (Hambro et al., 2022b). This model utilizes an LSTM architecture that incorporates representations from three encoders, which take observations as inputs. The LSTM network’s output is then fed into two separate heads: a policy head and a baseline head. The model architecture used both in online and offline settings consists of a joint backbone for both actor and critic. It takes as an input three components: main observation of the dungeon screen, blstats, and message. blstats refers to the player’s status information, such as health and hunger, and message refers to the textual information displayed to the player, such as notifications and warnings. blstats and message are processed using two layer MLP. The main observation of the dungeon screen is processed by embedding each character and color in an embedding lookup table which is later put into a grid processed by ResNet. For more details about Main screen encoder refer to (Tuyls et al., 2023). The components are encoded, and are merged before passing to LSTM. This baseline allows for fast training but struggles with learning complex behaviours required for certain roles in the game. More details about the architecture can be found in (Tuyls et al., 2023; Petrenko et al., 2020). The model hyperparameters are shown in Table 1 – analogical to Table 6 from (Petrenko et al., 2020).
@@
The knowledge retention methods presented in this paper use a subset of the NetHack Learning Dataset (NLD) collected by (Hambro et al., 2022b) called NLD-AA. It contains over 3 billion state-action-score transitions and metadata from 100,000 games collected from the winning bot of the NetHack Challenge (Hambro et al., 2022a). In particular, we use about 8000 games of Human Monk. This character was chosen because it was extensively evaluated in the previous work (Hambro et al., 2022b) and because the game setup for the Human Monk is relatively straightforward, as it does not require the agent to manage the inventory. The bot is based on the ’AutoAscend’ team solution, a symbolic agent that leverages human knowledge and hand-crafted heuristics to progress in the game. Its decision-making module is based on a behavior tree model.
@@
Table 1: Hyperparameters of the model used in NLE. For the most part, we use hyperparameters values from (Hambro et al., 2022b).
@@ C.2 Distillation-based methods
$\mathcal{L}_{BC}(\theta)=\mathbb{E}_{s\sim\mathcal{B}}[D_{KL}^{s}(\pi_{\theta}\;\|\;\pi_{*})],$ (3)
@@ Appendix D Additional NetHack results
DQN-Offline (Hambro et al., 2022b) 0.0 ± 0.0
CQL (Hambro et al., 2022b) 366 ± 35
IQL (Hambro et al., 2022b) 267 ± 28
BC (CDGPT5) (Hambro et al., 2022b, a) 1059 ± 159
@@
From Scratch + KS (Hambro et al., 2022b) 2090 ± 123
From Scratch + BC (Hambro et al., 2022b) 2809 ± 103

### v3

@@ 2 Forgetting of pre-trained capabilities
In state coverage gap, we consider a pre-trained agent that is performing well mostly on Far and does not know how to behave on Close. However, when fine-tuned on Close, its behavior on Far will deteriorate considerably due to forgetting [^1] For a more thorough discussion on the nature of interference in RL we refer the reader to Schaul et al. (2019) and will have to be re-acquired. This setting is representative of common transfer RL scenarios (Parisotto et al., 2015; Rusu et al., 2016, 2022), see also the top row of Figure 1 and Figure 2 for illustration.
@@ 3 Experimental setup
NetHack Learning Environment (Küttler et al., 2020) is a complex game consisting of procedurally generated multi-level dungeons. Since their layouts are randomly generated in each run, the player has to learn a general strategy rather than memorize solutions. NetHack is stochastic and requires mastering diverse skills, such as maze navigation, searching for food, fighting, and casting spells. It has been a popular video game for decades that recently has become a challenging testbed at the forefront of RL research (Hambro et al., 2022a; Piterbarg et al., 2023; Klissarov et al., 2023). Due to computational constraints, we focus solely on a single setting in our experiments, i.e., Human Monk. The code is available at https://github.com/BartekCupial/finetuning-RL-as-CL.
@@ 5 Analysis: forgetting of pre-trained capabilities hinders RL fine-tuning
We study the extent to which knowledge retention techniques mitigate the negative effects of imperfect cloning gap on two levels representing Far states: level $4$ and Sokoban level [^2] In NetHack, the Sokoban level is a branch of the dungeon modeled and named after an NP-hard game where the goal is to push boxes on target locations, see NetHack wiki and Appendix B.1., see Figure 5. The performance of fine-tuning on level 4 can be temporarily enhanced by EWC and consistently improved by KS or BC, which is in line with the results presented in Figure 3(a). Solving the Sokoban level does not yield immediate rewards so the vanilla fine-tuning agent pursues other strategies that are more beneficial in the short term. As such, it is not surprising that this particular behavior is forgotten. However, forgetting this skill will be hurtful in the long term, since completing the Sokoban levels unlocks a variety of equipment that is crucial for high performance during the rest of the game. Differentiating between behaviors that should be forgotten and those that should be kept is an important future direction for knowledge retention methods.
The Sokoban results allow us to get some insights into the qualitative differences between the KS and BC. Namely, KS struggles with sustaining the performance on Sokoban, as uses trajectories gathered by the online policy. These do not contain any Sokoban levels at the start of the fine-tuning, as Sokoban is only encountered in the deeper parts of the dungeon. Conversely, BC uses data gathered by the expert and, as a result, constantly rehearses the correct way of solving this puzzle. As such, we note that both BC and KS have their specific advantages. We identify designing methods that combine these improvements as important future work. See Appendix D for additional NetHack metrics.
@@
Other considerations Additionally, we note that choosing the most appropriate knowledge retention method for the problem at hand may depend on other constraints. In particular, if no prior data can be utilized during the fine-tuning, then BC and EM cannot be used, and one should instead leverage EWC, or apply Kickstarting that distills the knowledge on the online data. On the other hand, if one wishes to minimize computational complexity, EWC might be the best choice, as the other methods require processing more examples per training step. Finally, if there are restrictions on memory, one should weigh the cost of keeping the data (needed for BC, EM) against the cost of keeping the parameters of the pre-trained model (needed for EWC, KS).
@@ 6 Related Work
Offline to Online Reinforcement Learning
Recent work explored techniques for efficiently transitioning from offline to online reinforcement learning. Ball et al. (2023) use symmetric sampling of offline and online data and combine it with layer normalization and ensembles in an off-policy setting. Lee et al. (2022c) propose using a network for measuring “online-ness” of data and prioritizing samples in a replay buffer according to that measure. Nakamoto et al. (2024) modify Conservative Q-Learning to train on a mixture of the offline data and the new online data, weighted in some proportion during fine-tuning. We highlight that mixing new data with old data can be viewed as a knowledge retention technique similar to Episodic Memory. Although these approaches are relevant to our study and we see testing them as important future work, we use behavioral cloning in pre-training for simplicity, especially as it has been shown to outperform offline RL methods in the NetHack domain (Hambro et al., 2022b).
Impact of interdependence between Far and Close
The relation between Far and Close states has an important impact on the degree of forgetting, which might be understood through the lens of CL literature on task similarity. For example, Lee et al. (2021) find that intermediate task similarity levels lead to the highest degrees of forgetting. Evron et al. (2022) reach a similar conclusion in the linear regression setting when a given task is seen only once, but also find that high similarity causes most forgetting when repeatedly revisiting tasks. Furthermore, Evron et al. (2024) suggest that this behavior might be explained by heavy overparameterization, since in non-overparameterized cases forgetting grows monotonically as the task difference increases.
Generalization to multi-task setting
While our work focuses on single-task fine-tuning, prior research has explored fine-tuning on multiple unseen tasks. Yang et al. (2023) compared offline RL methods with imitation learning in a 2D goal-reaching aiming to test generalization to unseen goals. Mandi et al. (2022) showed that multi-task pre-training with fine-tuning often outperformed meta-reinforcement learning approaches in adaptation tasks with high task diversity and strictly unseen test tasks. At the same time, we believe studying single-task fine-tuning in NetHack provides valuable preliminary insights into this problem, as the game’s procedural generation on each run requires flexibly applying learned skills to adapt to new contexts.
@@ 7 Limitations & Conclusions
While our study focuses on the RL setting, some of its findings might have a broader scope. Non-stationary dynamics might also emerge in the supervised learning i.i.d. setting when the model sequentially acquires increasingly sophisticated skills (e.g., LLMs first learn simple grammar and understand advanced skills only much later) (Evanson et al., 2023; Luo et al., 2023). This suggests that the principles of knowledge retention and forgetting we explored could be relevant beyond the specific RL scenarios we tested, potentially impacting a wide range of learning systems that evolve over time. A comprehensive examination of these dynamics across different learning models and environments remains a crucial area for future research.
Impact statement
@@ B.1 NetHack
We fine-tune the model pre-trained by Tuyls et al. (2023), which scales up (from 6M to 33M parameters) and modifies the solution proposed by the ’Chaotic Dwarven GPT-5’ team, which is based on Sample Factory (Petrenko et al., 2020) that was also used in (Hambro et al., 2022c). This model utilizes an LSTM architecture that incorporates representations from three encoders, which take observations as inputs. The LSTM network’s output is then fed into two separate heads: a policy head and a baseline head. The model architecture used both in online and offline settings consists of a joint backbone for both actor and critic. It takes as an input three components: main observation of the dungeon screen, blstats, and message. blstats refers to the player’s status information, such as health and hunger, and message refers to the textual information displayed to the player, such as notifications and warnings. blstats and message are processed using two layer MLP. The main observation of the dungeon screen is processed by embedding each character and color in an embedding lookup table which is later put into a grid processed by ResNet. For more details about Main screen encoder refer to (Tuyls et al., 2023). The components are encoded, and are merged before passing to LSTM. This baseline allows for fast training but struggles with learning complex behaviours required for certain roles in the game. More details about the architecture can be found in (Tuyls et al., 2023; Petrenko et al., 2020). The model hyperparameters are shown in Table 1 – analogical to Table 6 from (Petrenko et al., 2020).
@@
The knowledge retention methods presented in this paper use a subset of the NetHack Learning Dataset (NLD) collected by (Hambro et al., 2022c) called NLD-AA. It contains over 3 billion state-action-score transitions and metadata from 100,000 games collected from the winning bot of the NetHack Challenge (Hambro et al., 2022a). In particular, we use about 8000 games of Human Monk. This character was chosen because it was extensively evaluated in the previous work (Hambro et al., 2022c) and because the game setup for the Human Monk is relatively straightforward, as it does not require the agent to manage the inventory. The bot is based on the ’AutoAscend’ team solution, a symbolic agent that leverages human knowledge and hand-crafted heuristics to progress in the game. Its decision-making module is based on a behavior tree model.
@@
Table 1: Hyperparameters of the model used in NLE. For the most part, we use hyperparameters values from (Hambro et al., 2022c).
@@ C.2 Distillation-based methods
$\mathcal{L}_{BC}(\theta)=\mathbb{E}_{s\sim\mathcal{B}}[D_{KL}^{s}(\pi_{*}(\cdot|s)\;\|\;\pi_{\theta}(\cdot|s))],$ (3)
@@ Appendix D Additional NetHack results
DQN-Offline (Hambro et al., 2022c) 0.0 ± 0.0
CQL (Hambro et al., 2022c) 366 ± 35
IQL (Hambro et al., 2022c) 267 ± 28
BC (CDGPT5) (Hambro et al., 2022c, a) 1059 ± 159
@@
From Scratch + KS (Hambro et al., 2022c) 2090 ± 123
From Scratch + BC (Hambro et al., 2022c) 2809 ± 103
@@ Appendix F Analysis of forgetting in robotic manipulation tasks
Figure 22: The performance on a robotic sequence where the sequence consists of the same tasks, but with observations translated by a constant $c$. We can observe forgetting even for small perturbations ($c=0.1$).
Impact of task difference
The severity of forgetting is deeply connected to how different Far and Close tasks are to each other. We refer the reader to Section 6 for a short description of prior continual learning papers on this problem, and here we perform a simple experiment on this issue. We construct a RoboticSequence consisting of tasks peg unplug (translated), push wall (translated), peg unplug, push wall and use a model pre-trained on the last two tasks. (Translated) means that the observation vectors are shifted by a constant $c$. This is a very basic form of state perturbation. In this case, the non-translated (translated resp.) stages correspond to Far (Close resp.) states. We run vanilla fine-tuning experiments with values of $c\in(0.01,0.1,1,10)$. We observe no forgetting for $c=0.01$, partial forgetting for $c=0.1$, and total forgetting for $c=1$, and $c=10$. We treat this result as initial evidence supporting the claim that even small discrepancies between far and close states might lead to forgetting.